Shopping Cart
Your Cart is Empty
There was an error with PayPalClick here to try again
CelebrateThank you for your business!You should be receiving an order confirmation from Paypal shortly.Exit Shopping Cart

25/09/2019 - Midweek progress from Jeff -

"A day at Kidderminster TMD today to continue fitting / wiring of "A" cab desk components to D1013."


22/09/2019 - Update pictures from Sunday's preparation work in readiness for the equalising beam and pedestal returning this week.

07/09/2019 - Update pictures of the weekend's work from Russ, Darren, Lewis and Chris

07/09/2019 - Update of the mid-week and weekend's work

24/08/2019 - And here we go again..... same bogie defect as 2018 below..... we have more experience at this now, and the team worked hard over the last two weekends to lift D1062 and remove the "B" End bogie for identical repairs to a tilted axle.

15/05/18 to 18/05/18 - And It's Gala time!...... And the work still goes on!

12/05/18 - Another Bogie Weekend...and preparations as arrivals for the Gala appear in shot...

14/04/2018 - Bogies Again!

31/10/15 - A Halloween Day To Remember!

They say the sun does shine on the righteous!.......... And also, it seems, on our working members (!) who enjoyed a well earned breather in the sunshine earlier today. Lots done on several fronts - lets take a closer look.....

Left to Right: Paul, Jeff, Roger, Russ, Stuart, Lyn, Jack, Mark and Rob.

The key work on the spare bogie continued today with various components being encouraged to leave their state of encasement in favour of refurbishment and a future return to a wonderful working life!

Now those overalls are just far to clean to be doing that job. Peering deep into the innards, the work continues to challenge our illustrious team who get to grips with the heart of the bogie in the Autumn sunshire.

Pete Sandham takes a turn between the bogie frames. Russ, Robin and Paul pondering the next move.

Injector tests have produced 4 fails out of six (so far) on the injectors removed from D1062 B End this weekend. Nothing we can't put right! 

Left: Meanwhile, work continues with the pre-heater control panel rewiring - Roger hard at work. 

Right: Fruits of their labours - injectors safely removed from B end ready for inspection, calibration tests and refit. Paul and Jack Walker do the honours.

Injector on the test rig earlier today. This one was one of the two to "pass".